OpenStudy (anonymous):

In the answer space below, provide the larger of the two positive integers that add up to 60 and have the largest possible product.

OpenStudy (radar):

x + y = 60 x = 60 - y Now solve for xy (60 - y) y = 60y - y^2 Call product P so then P = 60y - y^2. Now differentiate P' = -2y + 60 For Maximum set P' to 0 0 + -2y + 60 now solve for y -2y = -60 y=30 now solve for x x + 30 = 60 x=60-30 = 30 the answer is that was it.

OpenStudy (anonymous):

thank you